A particle start moving with a constant angular acceleration Alpha in a circular path .The time at which the magnitude of tangential and radial acceleration are equal is:

Grade:11

1 Answers

Arun
5 years ago
As we realize that 
Centripetal speeding up = v^2/r 
To start with condition of movement 
ω = ωo+ αt 
As beginning rakish speed is 0 
ω = αt 
v/r = at 
v = atr. 
Extraneous increasing speed = centripetal quickening 
αr = a 
ar = v^2/r 
ar = (atr)2/r 
a =  a^2t^2 
t = 1/sqrt (a)